what is the largest nunber A such that the graphs of :
x^2 = y^2 and (x-A)^2 +y^2 =1 intersect?

Answer by Alan3354(69443)

what is the largest nunber A such that the graphs of :
x^2 = y^2 and (x-A)^2 +y^2 =1 intersect?
-----------
x^2 = y^2 is a degenerate (I think that's the right term) hyperbola. It's 2 straight lines with slopes of -1 and +1 thru the Origin.
----
(x-A)^2 +y^2 = 1 is a circle of radius 1, center at (A,0) - on the x-axis.
===================
find the 2 tangent points, then find A.
----
email via the TY note for help, or to check you work.
PS Graphing is helpful.